A 9.7-g bullet is fired into a stationary block of wood having mass m = 5.02 kg. The bullet imbeds into the block. The speed of the bullet-plus-wood combination immediately after the collision is 0.610 m/s. What was the original speed of the bullet? (Express your answer with four significant figures.)
Answer #1

Given :

Mass of bullet = M​​​​​​1 = 9.7 * 10-3 kg

Mass of block of wood = m = 5.02 kg

Let initial speed of bullet = V​​​​​​1 m/s

Final velocity of bullet + wood = 0.610 m/s

Now applying conservation of momentum :-

Initial momentum of bullet + wood = finlal momentum of bullet+ wood

M​​​​​​1* V​​​​​​1 + m* v = V​​​​​​f ( M​​​​​​1 + m)  

9.7 * 10-3 * V​​​​​​1 + 0 = 0.610 * ( 9.7 * 10-3 + 5.02 ) ( v= 0 , as

wood is at rest)

V​​​​​​1 = 316.30 m/s

Original speed of bullet = V​​​​​​1= 316.30 m/s